What is the molar concertation of 10g of salt (NaCl) in 100 mL of water?
1
Expert's answer
2020-10-30T06:40:55-0400

Molar concentration =(mass/molar mass) ×(1000/volume in mL)

Molar mass of NaCl = 23+35.5= 58.5g/mol

Molar concentration = (10/58.5)×(1000/100)

Molar concentration = 1.709M